# Quantifying individual and team performance in cricket

### v1

@@ III Discussion
Our results might be subject to certain limitations. Firstly, the volume of One Day International (ODI) matches and the resulting dataset size are not fully extensive. Future research could benefit from incorporating data from diverse cricket formats such as Test matches, T20 games, and franchise leagues like the IPL, which might yield deeper insights into varied player performance patterns. Supplementing our analysis with data from various professional levels may reveal further patterns in players’ mobility and skill levels. Additionally, our dataset focuses on end-match statistics, omitting the nuanced temporal dynamics within individual games. While the final scorecard offers substantial information, some facets of performance may only become apparent with a more granular data analysis.
All in all, our work reveals intriguing patterns of individual and team performance in cricket. We hope that our analysis motivates further research into synergistic individual efforts and team dynamics.
IV Methods
IV.1 Data
In our study, we extracted data on 4566 One Day International (ODI) matches in men’s cricket, involving 2882 unique players, sourced from howstat.com. This dataset spans from 1971 to 2023. For each match, we extracted information including match-date, teams, runs scored, wickets taken, and overs played by each team, along with player names and their contributions in terms of batting and bowling. This included details like batting position, number of balls faced, number of fours and sixes hit, number of overs bowled, number of runs conceded, number of maiden overs, and number of wickets taken. Furthermore, data on the captain and wicketkeeper for each match, as well as the winner of the match, were also collected.
IV.2 Cricket Terms
IV.2.1 Gameplay
A One Day International cricket game consists of each team playing 50 overs, with 11 players per side and one over comprising 6 balls  [^2] The number of overs per side in a match was originally 60 before all teams adopted the 50-over format. However, the number of such matches is low, so we do not account for this factor.. The playing field is circular, featuring a 20.11-meter pitch at its centre, with stumps or wickets at both ends. The gameplay is similar to baseball with few key differences and some terminology changes. The bowler or the thrower projects the ball towards the batsman or hitter, aiming for it to bounce once before reaching them. The batsman’s goal is to hit the ball hard so that it crosses the boundary for 4 or 6 runs. If the bowl doesn’t cross the boundary, the batsmen run between the wickets for additional runs until the players from the other team collect the ball and throw it back towards the wickets. The bowler’s objective is to hit the stumps or induce the batsman to hit the ball such that a fielder can catch it before it touches the ground, resulting in the batsman being declared out. The innings of a team batting first concludes when either 50 overs are bowled or 10 batsmen are declared out. The aim for the second team is to score at least 1 run more than the first team. The match ends when the second team scores more runs than the first team in 50 overs or the second team fails to reach the target either due to all players declared out or 50 overs being done. The team with a higher score is declared as the winner.
IV.2.2 Type of Players
Batsman: We consider players as batsmen who have played at least 25 matches and batted at position 7 or above in at least 50% of their matches. We have 607 batsmen in our dataset after applying this criteria. Like in baseball, the job of the batsman is to score as many runs as possible by hitting the ball out of bounds or by running between the wickets if the ball remains within bounds.
Bowlers: We classify players as bowlers who have played at least 25 matches and bowled in at least 50% of the matches they played. We have 594 bowlers in our dataset after applying this criterion. The main task of the bowlers is to dislodge the stumps or force the batsman to hit the ball towards the fielder.
All-rounders: We consider all-rounders as players who have played at least 25 matches and are classified as both batsmen and bowlers using the above definitions.
Captains: A captain for a match is the leader of the team who directs the team and makes strategic in-game decisions. In our analysis, we consider players as captains if they have captained the team in at least 50 matches.
Openers and Non-openers: We characterize openers as players who have batted at first or second position at least 10 times in their career, while non-openers as players who have batted at positions three to six at least 10 times in their career.
Wicket-keepers: The dataset contains an attribute for each player specifying if they were the wicket-keepers in the match. The wicket-keepers ensure that the bowl doesn’t go beyond them in case the batsman fails to hit the ball. They also try to dislodge the stumps if the batsman fails to make it to the wickets while scoring runs.
Fielders: All players on the bowling side who are not bowling or wicket-keeping are classified as fielder in our analysis. There are two main responsibilities of the fielder. First, they have to catch a ball hit by the batsman to dismiss them before the ball touches the ground. Second, in case the first scenario is not possible, the fielder has to collect the ball and throw it at either wickets to catch the batsman short of completing the run (thus, dismissing them) or stopping further runs.
IV.3 Null Models
IV.3.1 Best individual performance
In fig (1), we characterize the existence of random impact rule and hot streaks. In order to establish a baseline for these measures, we shuffle the timestamps of individual performances for each player 5000 times keeping the actual values of performance the same. In this way, we are breaking the temporal correlation between different performances. For each shuffling, we find the time stamp for the best ($N^{*}$), second best ($N^{**}$), and third best ($N^{***}$) performances of a player. We plot the PDF of this randomly shuffled data to compare the results from our dataset.
IV.3.2 Hot and cold streaks for teams
Figure (5) (a) showcases the existence of hot and cold streaks for teams. We consider all the teams that have played at least 200 matches with a definite result. By keeping the number of wins constant, we shuffle the timestamps of the match results $10^{4}$ times for each team. For each reshuffling, we compute the number of consecutive wins (hot) and losses (cold) denoted by $L_{streak}$. By taking the average across $10^{4}$ reshufflings, we compute $N_{rand}$ to highlight the significance of $N_{data}$ – the number of streaks of length $l$ or higher in the actual dataset.
@@ Acknowledgements
The authors warmly thank Aanjaneya Kumar for his initial involvement in the project and discussions of preliminary results. F.B. acknowledges support from the Air Force Office of Scientific Research under award number FA8655-22-1-7025.
Competing interests
The authors declare that they have no competing interests.
Additional information
@@
Correspondence should be addressed to BattistonF@ceu.edu
Supplemental material for “Quantifying individual and team performance in cricket”
@@
Figure S1: Inflation of performance across decades. We observe a 40% increase in the total team score from the 1980s to the 2020s. To ensure a fair comparison for players from different eras we normalize all the scores such that the average team score over the years is constant.
SM2 Distribution of inter-event match difference
@@
SM3 Correlators of career performance
Figure S3: Correlation of mean performance over different time periods to the full career performance for (top row) batsmen and (bottom row) bowlers. We notice that (a),(d) random, (b), (e) mid, and (c),(f) late career performances are correlated ($R^{2}>0.6$) to the full career performances giving evidence that intrinsic talent influences player performance.
SM4 Length of leadership
Figure S4: Distribution of leadership duration for captains who have led their team for at least 50 matches for batsmen and bowler captains. On average the batsmen lead their team in 215 matches, while the bowlers lead their team in approximately 240 matches before stepping down.
SM5 Effect of batting position on runs scored and balls faced.
@@
SM6 Contribution of non-specialists to team
@@
SM7 Individual contributions to team success

### v2

@@ IV Discussion
Our results might be subject to certain limitations. Cricket is a multi-faceted game, and while our analysis focuses on a wide range of indicators, it still does not consider all aspects of the game such as home advantage and differences among spinners and fast bowlers. Furthermore, the volume of One Day International (ODI) matches, and the resulting dataset size is not fully extensive. Future research could benefit from incorporating data from diverse cricket formats such as Test matches, T20 games, and franchise leagues like the IPL, which might yield deeper insights into varied player performance patterns. Supplementing our analysis with data from various professional levels may reveal further patterns in players’ mobility and skill levels. Additionally, our dataset focuses on end-match statistics, omitting the nuanced temporal dynamics within individual games. While the final scorecard offers substantial information, some facets of performance may only become apparent with a more granular data analysis.
All in all, our work reveals intriguing patterns of individual and team performance in One Day International cricket. We believe that our methodologies developed here could be readily applied without substantial modifications to other formats of the game such as T20 and Test cricket, as well as other team sports. Particularly suitable are teams with high specialisation such as baseball, American football, and volleyball. A comparative analysis of our results with a broader literature on sports can improve our understanding of human behaviour. We hope that our analysis motivates further research into synergistic individual efforts, in sports and more broadly in team dynamics.
@@ Acknowledgements
The authors thank Aanjaneya Kumar and Lorenzo Betti for the discussions which led to an improvement in the manuscript. F.B. acknowledges support from the Air Force Office of Scientific Research under award number FA8655-22-1-7025.
@@
Correspondence should be addressed to sadekar_onkar@phd.ceu.edu or battistonf@ceu.edu.
Supplemental material for “Individual and team performance in cricket”
@@
Figure S1: Inflation of performance across decades. We observe a 40% increase in the total team score from the 1980s to the 2020s. To ensure a fair comparison for players from different eras we normalise all the scores such that the average team score over the years is constant.
SM2 Presence of individual hot streaks
Figure S2: Presence of hot streaks for players for 4 different values for the number of bins of $\Delta N/N$. The ratio of original data ($n^{data}_{\Delta N}/n^{data}_{N}$) and randomised data ($n^{rnd}_{\Delta N}/n^{rnd}_{N}$) is always greater than 1 for the smallest bins, while it is always smaller than 1 for the largest bins indicating that the top performances tend to be more clustered than expected at random. All the results are statistically significant based on the Mann-Whitney U test.
SM3 Distribution of inter-event match difference
@@
SM4 Correlators of career performance
Figure S4: Correlation of mean performance over different time periods to the full career performance for (top row) batsmen and (bottom row) bowlers. We notice that (a),(d) random, (b), (e) mid, and (c),(f) late career performances are correlated ($R^{2}\sim 0.65$) to the full career performances giving evidence that intrinsic talent influences player performance.
SM5 Length of leadership
Figure S5: Distribution of leadership duration for captains who have led their team for at least 15 matches for batsmen and bowler captains. On average the batsmen lead their team in 211 matches, while the bowlers lead their team in approximately 232 matches before stepping down.
SM6 Bowler-captains’ burden
Figure S6: Boxplot showing the distribution of fraction of overs bowled by the captain (bottom) and the bowler who bowls the maximum number of overs in a given match (top). Only 31% of times the captain bowled the maximum number of overs. The average fraction bowled by the most active bowler is 0.23, while the captain bowler bowls on average 0.14 fraction of total overs. The Mann-Whitney U test to distinguish the two distributions yields a p value $<0.001$.
SM7 Effect of batting position on runs scored and balls faced.
@@
SM8 Contribution of non-specialists to team
@@
SM9 Individual contributions to team success